• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 16:19:39 on November23,2009

    Entry number 302100

    keyword=Day shift summary

    11/23/2009 Day shift summary

    Shift workers: J. P. Chen (SL), R. Michaels (TO), C. Jen (3rd)

    8:00 production with 110 uA on 20 cm LD2 (loop1).
    8:15 disk 2 (adaql1/data2) nearly full. Bob changed it to disk 4
    (adaql1/data4) (Log entry 302029).
    9:40 coda crash
    10:30 coda crash
    12:05 stoppde production. started ramping down right dipole to get ready for polarity change.
    Continue run while waiting for BCM calibration to start.
    13:11 BCM calibration and boiling tests.
    14:15 requested controlled access for polarity change
    15:30 Heidi/Todd completed polarity switch. Jack is still working on the left Q1.
    Started to set the spectrometer to R=2.63 GeV, L=3.66 GeV.
    Cycled Q2/Q3 right to 1400 A
    cycled left Q2 to 1600 and Q3 to 1400 (per Jack's instruction).
    leave Q1 alone for now.
    15:45 All magnets set except right dipole (still ramping up). Jack fixed left Q1.
    A question on right Q3, which did not switch sign (gauss meter reading) aftre polarity switch. Still investigating at the end of shift, while waiting for right dipole ramping up.

    run information:

    Happex runs
    14168 8:05-8:15, 13K, 110 uA on LD2, short run due to disk changed from /data2 to /data4 (log entry 302029).
    14169 8:18-9:13, 100K
    14170 9:20-9:40, 34K, Coda crashed
    14171 9:45-10:30, 34K, coda crashed again
    14172 10:44-11:44, 100K
    14173 11:44-12:05, 37K, stopped for ramping down right dipole for polarity change.
    14174 12:10-12:58 right dipole ramping down.
    14175 13:11-13:55, BCM calibration
    14176 boiling test

    HRS_L/R runs
    5257/25706 7:05-8:31/8:41, 5.7M/6M, 110 uA on LD2,
    5258/25707 8:35/8:45-10:01-10:24, 5.7M/6M
    5259/25708 10:04/10:28-11:27/11:31, 5.7M/4M
    5260/25709 11:28/11:33-12:06, 2.6M/2.1M.
    5261/25710 12:20/12:12-12:58, Right dipole ramping down.
    5262/25711 BCM calibration
    5263/25712 boiling test